OpenStudy (anonymous):

Do me a solid, Will medal (mixed fractions) 5 1/6- 2 2/3

OpenStudy (raffle_snaffle):

and what is 2*(2/3) = ?

OpenStudy (raffle_snaffle):

5*(1/6) = ?

OpenStudy (abdullahm):

2 2/3 is not the same as 2*(2/3)

OpenStudy (anonymous):

yea^

OpenStudy (raffle_snaffle):

okay is it 22/3 ?

OpenStudy (anonymous):

no. it says they're mixed fractions bruh

OpenStudy (raffle_snaffle):

I was asking her to multiply 2 by the numerator

OpenStudy (abdullahm):

@sheyennepowell_ First convert the fractions into improper fractions: Hint: \(\bf\Large a\frac{b}{c} = \frac{(a\times c)+b}{c}\)

OpenStudy (anonymous):

yep

OpenStudy (abdullahm):

Do that for both fractions and let me know what you get. :)

OpenStudy (anonymous):

31/6 and 17/3

OpenStudy (abdullahm):

Try again for 2 2/3. The other one is correct. :)

OpenStudy (anonymous):

ohhh that's where I went wrong that's the only reason I asked on here

OpenStudy (anonymous):

thank you

OpenStudy (abdullahm):

It was my pleasure! :)

OpenStudy (anonymous):

medal me?